OpenStudy (anonymous):

How do I factor 8y^3-1000z^6?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

I got to 8(y^3-125z^6) but I don't know where to go from there.

OpenStudy (anonymous):

8(y-5z^2)(y^2+5yz^2+25z^425z^4)

OpenStudy (anonymous):

8(y^3 -125z^6) = 8(y-5z^2)(y^2 + 25z^4 + 5yz^2)

OpenStudy (anonymous):

Yes but how did you get that? I'm confused on the factoring part.

OpenStudy (anonymous):

the second bracket can be again factorized... do it usself its quadratic in y

OpenStudy (anonymous):

u mean factor out y^3?

OpenStudy (anonymous):

\[a ^{3} - b ^{3} = \left( a-b \right)\left( a ^{2} + b ^{2} +ab \right)\]

OpenStudy (anonymous):

@ineedhelp911 got it!!

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

this is difficult. where did the 5zy^2 come from?

OpenStudy (anonymous):

ab part u multiply the 2 things ...

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

Difference of cubes \[8y^3-1000z^6=(2y)^3-(10z^2)^3\]